# Subsampled Ensemble Can Improve Generalization Tail Exponentially

### v3

@@ C.1 Preliminaries
This completes the proof. $\Box$
Next, we present our sharper concentration bound for U-statistics with binary kernels:
Lemma 2 (Concentration bound for U-statistics with binary kernels)
Proof of Lemma 2. We first consider the direction $U-\mathbb{E}\left[\kappa\right]\geq\epsilon$. Let
@@
where we apply the Markov inequality in $(i)$, step $(ii)$ is due to Lemma 1, and step $(iii)$ uses Jensen’s inequality and the convexity of the exponential function. Due to independence, $\tilde{\kappa}$ can be viewed as the sample average of $\hat{n}$ i.i.d. Bernoulli random variables, i.e., $\tilde{\kappa}\sim\frac{1}{\hat{n}}\sum_{i=1}^{\hat{n}}\operatorname{Bernoulli}\left(\mathbb{E}\left[\kappa\right]\right)$. Hence, we have that
@@
Since $D_{\operatorname{KL}}(p\|q)=D_{\operatorname{KL}}(1-p\|1-q)$, the bound for the reverse side $U-\mathbb{E}\left[\kappa\right]\leq-\epsilon$ then follows by applying the first bound to the flipped binary kernel $1-\kappa$ and $1-U$. This completes the proof of Lemma 2. $\Box$
Next lemma gives lower bounds for KL divergences which help analyze the bounds in Lemma 2:
Lemma 3
$D_{\operatorname{KL}}(p\|q)\geq p\ln\frac{p}{q}+q-p.$ (17)
Proof of Lemma 3. To show (17), some basic calculus shows that for any fixed $q$, the function $g(p):=(1-p)\ln\frac{1-p}{1-q}$ is convex in $p$, and we have that
@@
This completes the proof of Lemma 3. $\Box$
$\mathbb{A}(z_{1},\ldots,z_{k};\omega):\mathcal{Z}^{k}\times\Omega\to 2^{\Theta}$ (19)
where $\omega$ denotes algorithmic randomness that is independent of the data $\left\{z_{1},\ldots,z_{k}\right\}$. Each of our proposed algorithms attempts to solve the probability-maximization problem
The problem actually solved with a finite number of subsamples is
Specifically, Algorithm 1 uses
$\mathbb{A}(z^{*}_{1},\ldots,z^{*}_{k};\omega)=\left\{\mathcal{A}(z^{*}_{1},\ldots,z^{*}_{k};\omega)\right\}$ (23)
$\mathbb{A}(z^{*}_{1},\ldots,z^{*}_{k_{2}};\omega)=\left\{\theta\in\mathcal{S}:\frac{1}{k_{2}}\sum_{i=1}^{k_{2}}l(\theta,z^{*}_{i})\leq\min_{\theta^{\prime}\in\mathcal{S}}\frac{1}{k_{2}}\sum_{i=1}^{k_{2}}l(\theta^{\prime},z^{*}_{i})+\epsilon\right\}$ (24)
conditioned on the solution set $\mathcal{S}$ retrieved in Phase I. Note that no algorithmic randomness is involved in (24) once the set $\mathcal{S}$ is given. We define:
Definition 2
For any $\delta\in[0,1]$, let
be the set of $\delta$-optimal solutions of problem (21). Let
and
Definition 3
Let
C.2 Proof for Theorem 1
We consider Algorithm 3, a more general version of Algorithm 1 that operates on the set-valued learning algorithm $\mathbb{A}$ in (19) and reduces to exactly Algorithm 1 in the special case (23). Again we omit the algorithmic randomness $\omega$ in $\mathbb{A}$ for convenience.
2:  for $b=1$ to $B$  do
4:  end for
We have the following finite-sample result for Algorithm 3:
Theorem 3 (Finite-sample bound for Algorithm 3)
Consider discrete decision space $\Theta$. Recall $p_{k}(\theta)$ defined in (21). Let $p_{k}^{\max}:=\max_{\theta\in\Theta}p_{k}(\theta)$ and
where $\max_{\theta\in\Theta\backslash\Theta^{\delta}}p_{k}(\theta)$ evaluates to 00 if $\Theta\backslash\Theta^{\delta}$ is empty. For every $k\leq n$ and $\delta\geq 0$ such that $\bar{\eta}_{k,\delta}>0$, the solution output by Algorithm 3 satisfies that
$\leq$ $\left|\Theta\right|\Bigg{[}\exp\left(-\dfrac{n}{2k}\cdot D_{\operatorname{KL}}\left(p_{k}^{\max}-\dfrac{3\eta}{4}\Big{\|}p_{k}^{\max}-\eta\right)\right)+2\exp\left(-\dfrac{n}{2k}\cdot D_{\operatorname{KL}}\left(p_{k}^{\max}-\dfrac{\eta}{4}\Big{\|}p_{k}^{\max}\right)\right)+$
$\exp\left(-\dfrac{B}{24}\cdot\dfrac{\eta^{2}}{\min\left(p_{k}^{\max},1-p_{k}^{\max}\right)+3\eta/4}\right)+$
$\mathbbm{1}\left(p_{k}^{\max}+\dfrac{\eta}{4}\leq 1\right)\cdot\exp\Big{(}-\dfrac{n}{2k}\cdot D_{\operatorname{KL}}\left(p_{k}^{\max}+\dfrac{\eta}{4}\Big{\|}p_{k}^{\max}\right)-\dfrac{B}{24}\cdot\dfrac{\eta^{2}}{1-p_{k}^{\max}+\eta/4}\Big{)}\Bigg{]}$
$\left|\Theta\right|\left(3\min\left(e^{-2/5},C_{1}\max(1-p_{k}^{\max},\;\max_{\theta\in\Theta\backslash\Theta^{\delta}}p_{k}(\theta))\right)^{\frac{n}{C_{2}k}}+\exp\left(-\frac{B}{C_{3}}\right)\right),$ (31)
Proof of Theorem 3. We first prove excess risk tail bounds for the problem (21), split into two lemmas, Lemmas 4 and 5 below.
Lemma 4
Consider discrete decision space $\Theta$. Recall from Definition 2 that $p_{k}^{\max}=p_{k}(\theta_{k}^{\max})$ holds for $\theta_{k}^{\max}$. For every $0\leq\epsilon\leq\delta\leq p_{k}^{\max}$, it holds that
Proof of Lemma 4. By Definition 2, we observe the following equivalence
@@ C.2 Proof for Theorem 1
$\mathbb{P}\left({\hat{p}}_{k}(\theta)\geq{\hat{p}}_{k}\left(\theta_{k}^{\max}\right)-\epsilon\right)$ (32)
$\leq$ ($\mathbb{P}\left(\left\{{\hat{p}}_{k}(\theta)\geq p_{k}(\theta_{k}^{\max})-\dfrac{\delta+\epsilon}{2}\right\}\cap\left\{{\hat{p}}_{k}\left(\theta_{k}^{\max}\right)\leq p_{k}(\theta_{k}^{\max})-\dfrac{\delta-\epsilon}{2}\right\}\right)$
@@
Note that, with $\kappa(z_{1},\ldots,z_{k};\omega):=\mathbf{1}\left(\theta\in\mathbb{A}(z_{1},\ldots,z_{k};\omega)\right)$, the probability ${\hat{p}}_{k}(\theta)$ can be viewed as a U-statistic with the kernel $\kappa^{*}(z_{1},\ldots,z_{k}):=\mathbb{E}\left[\kappa(z_{1},\ldots,z_{k};\omega)|z_{1},\ldots,z_{k}\right]$. A similar representation holds for ${\hat{p}}_{k}\left(\theta_{k}^{\max}\right)$ as well. Therefore, we can apply Lemma 2 to conclude that
@@
which completes the proof of Lemma 4. $\Box$
Lemma 5
$\mathbb{P}_{*}\left(\hat{\theta}_{n}\notin\widehat{\mathcal{P}}^{\epsilon}_{k}\right)\leq\left|\Theta\right|\cdot\exp\left(-\dfrac{B}{6}\cdot\dfrac{\epsilon^{2}}{\min\left(\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta_{k}^{\max})\right)+\epsilon}\right),$
[…]
Proof of Lemma 5. We observe that $\bar{p}_{k}(\theta)$ is an conditionally unbiased estimator for $\hat{p}_{k}(\theta)$, i.e., $\mathbb{E}_{*}\left[\bar{p}_{k}(\theta)\right]={\hat{p}}_{k}(\theta)$. We can express the difference between $\bar{p}_{k}(\theta)$ and $\bar{p}_{k}(\theta_{k}^{\max})$ as the sample average
@@
$\mathbb{P}_{*}\left(\bar{p}_{k}(\theta)-\bar{p}_{k}(\hat{\theta}_{k}^{\max})-(\hat{p}_{k}(\theta)-\hat{p}_{k}(\theta_{k}^{\max}))\geq t\right)$ (33)
$\leq$ $\exp\left(-B\cdot\dfrac{t^{2}}{2\mathrm{Var}_{*}(\mathbbm{1}_{\theta}^{*}-\mathbbm{1}_{\theta_{k}^{\max}}^{*})+4/3\cdot t}\right).$
@@
$\mathrm{Var}_{*}(\mathbbm{1}_{\theta}^{*}-\mathbbm{1}_{\theta_{k}^{\max}}^{*})$$\leq$ $\mathbb{E}_{*}\left[(\mathbbm{1}_{\theta}^{*}-\mathbbm{1}_{\theta_{k}^{\max}}^{*})^{2}\right]$
$\leq$ $\hat{p}_{k}(\theta)+\hat{p}_{k}(\theta_{k}^{\max})\leq 2\hat{p}_{k}(\theta_{k}^{\max}),$
@@
we have $\mathrm{Var}_{*}(\mathbbm{1}_{\theta}^{*}-\mathbbm{1}_{\theta_{k}^{\max}}^{*})\leq 2\min(\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta))$. Substituting this bound to (33) and taking $t=\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta)$ lead to
$\mathbb{P}_{*}\left(\bar{p}_{k}(\theta)-\bar{p}_{k}(\hat{\theta}_{k}^{\max})\geq 0\right)$$\leq$ $\exp\left(-B\cdot\dfrac{(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))^{2}}{4\min(\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta))+4/3\cdot(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))}\right)$
$\leq$ $\exp\left(-B\cdot\dfrac{(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))^{2}}{4\min(\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta_{k}^{\max}))+16/3\cdot(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))}\right)$
$\leq$ $\exp\left(-\dfrac{B}{6}\cdot\dfrac{(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))^{2}}{\min(\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta_{k}^{\max}))+\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta)}\right).$
@@
$\leq\sum_{\theta\in\Theta\backslash\widehat{\mathcal{P}}^{\epsilon}_{k}}\exp\left(-\dfrac{B}{6}\cdot\dfrac{(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))^{2}}{\min(\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta_{k}^{\max}))+\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta)}\right).$
@@
Note that the function $x^{2}/(\min(\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta_{k}^{\max}))+x)$ in $x\in[0,1]$ is monotonically increasing and that $\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta)>\epsilon$ for all $\theta\in\Theta\backslash\widehat{\mathcal{P}}^{\epsilon}_{k}$. Therefore, we can further bound the probability as
@@
$\mathbb{P}_{*}\left(\hat{\theta}_{n}\notin\widehat{\mathcal{P}}^{\epsilon}_{k}\right)$$\leq$ $\left|\Theta\backslash\widehat{\mathcal{P}}^{\epsilon}_{k}\right|\cdot\exp\left(-\dfrac{B}{6}\cdot\dfrac{\epsilon^{2}}{\min\left(\hat{p}_{k}^{\max},1-\hat{p}_{k}^{\max}\right)+\epsilon}\right).$
@@
Noting that $\left|\Theta\backslash\widehat{\mathcal{P}}^{\epsilon}_{k}\right|\leq\left|\Theta\right|$ completes the proof of Lemma 5. $\Box$
We are now ready for the proof of Theorem 3. We first note that, if $\bar{\eta}_{k,\delta}>0$, it follows from Definition 2 that
Therefore, for any $\eta\in(0,\;\bar{\eta}_{k,\delta})$, we can write that
$\mathbb{P}\left(\hat{\theta}_{n}\notin\Theta^{\delta}\right)\leq\mathbb{P}\left(\hat{\theta}_{n}\notin\mathcal{P}^{\eta}_{k}\right)$$\leq$ ($\mathbb{P}\left(\left\{\hat{\theta}_{n}\notin\widehat{\mathcal{P}}^{\eta/2}_{k}\right\}\cup\left\{\widehat{\mathcal{P}}^{\eta/2}_{k}\not\subseteq\mathcal{P}^{\eta}_{k}\right\}\right)$ (34)
$\leq$ $\mathbb{P}\left(\hat{\theta}_{n}\notin\widehat{\mathcal{P}}^{\eta/2}_{k}\right)+\mathbb{P}\left(\widehat{\mathcal{P}}^{\eta/2}_{k}\not\subseteq\mathcal{P}^{\eta}_{k}\right).$
We first evaluate the second probability on the right-hand side of (34). Lemma 4 gives that
$\mathbb{P}\left(\widehat{\mathcal{P}}^{\eta/2}_{k}\not\subseteq\mathcal{P}^{\eta}_{k}\right)\leq\left|\Theta\right|\Bigg{[}$ $\exp\left(-\dfrac{n}{2k}\cdot D_{\operatorname{KL}}\left(p_{k}^{\max}-\dfrac{3\eta}{4}\Big{\|}p_{k}^{\max}-\eta\right)\right)$ (35)
$+$ $\exp\left(-\dfrac{n}{2k}\cdot D_{\operatorname{KL}}\left(p_{k}^{\max}-\dfrac{\eta}{4}\Big{\|}p_{k}^{\max}\right)\right)\Bigg{]}.$
Next, by applying Lemma 5 with $\epsilon=\eta/2$, we can bound the first probability on the right-hand side of (34) as
$\mathbb{P}\left(\hat{\theta}_{n}\notin\widehat{\mathcal{P}}^{\eta/2}_{k}\right)\leq\left|\Theta\right|\cdot\mathbb{E}\left[\exp\left(-\dfrac{B}{24}\cdot\dfrac{\eta^{2}}{\min\left(\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta_{k}^{\max})\right)+\eta/2}\right)\right].$ (36)
$\mathbb{E}\left[\exp\left(-\dfrac{B}{24}\cdot\dfrac{\eta^{2}}{\min\left(\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta_{k}^{\max})\right)+\eta/2}\right)\right]$
$\mathbb{P}\left(\left|\hat{p}_{k}(\theta_{k}^{\max})-p_{k}^{\max}\right|<\frac{\eta}{4}\right)\cdot\exp\left(-\dfrac{B}{24}\cdot\dfrac{\eta^{2}}{\min\left(p_{k}^{\max},1-p_{k}^{\max}\right)+3\eta/4}\right)+$
$\leq$ $\mathbb{P}\left(\hat{p}_{k}(\theta_{k}^{\max})\leq p_{k}^{\max}-\frac{\eta}{4}\right)+\exp\left(-\dfrac{B}{24}\cdot\dfrac{\eta^{2}}{\min\left(p_{k}^{\max},1-p_{k}^{\max}\right)+3\eta/4}\right)+$
$\exp\left(-\dfrac{B}{24}\cdot\dfrac{\eta^{2}}{\min\left(p_{k}^{\max},1-p_{k}^{\max}\right)+3\eta/4}\right)+$
$\mathbbm{1}\left(p_{k}^{\max}+\dfrac{\eta}{4}\leq 1\right)\cdot\exp\left(-\dfrac{n}{2k}\cdot D_{\operatorname{KL}}\left(p_{k}^{\max}+\dfrac{\eta}{4}\Big{\|}p_{k}^{\max}\right)\right)\cdot\exp\left(-\dfrac{B}{24}\cdot\dfrac{\eta^{2}}{1-p_{k}^{\max}+\eta/4}\right)$
where inequality $(i)$ results from applying Lemma 2 with $\hat{p}_{k}(\theta_{k}^{\max})$, the U-statistic estimate for $p_{k}^{\max}$. Together, the above equations imply that
$\mathbb{P}\left(\hat{\theta}_{n}\notin\Theta^{\delta}\right)$
$\leq$ $\left|\Theta\right|\Bigg{[}\exp\left(-\dfrac{n}{2k}\cdot D_{\operatorname{KL}}\left(p_{k}^{\max}-\dfrac{3\eta}{4}\Big{\|}p_{k}^{\max}-\eta\right)\right)+$
$\exp\left(-\dfrac{B}{24}\cdot\dfrac{\eta^{2}}{\min\left(p_{k}^{\max},1-p_{k}^{\max}\right)+3\eta/4}\right)+$
and that $p_{k}^{\max}-\bar{\eta}_{k,\delta}\leq 1-\bar{\eta}_{k,\delta}\leq 1/5$, therefore by the bound (18) from Lemma 3, we can bound the first two terms as
$\exp\left(-\dfrac{n}{2k}\cdot D_{\operatorname{KL}}\left(p_{k}^{\max}-\dfrac{3\bar{\eta}_{k,\delta}}{4}\Big{\|}p_{k}^{\max}-\bar{\eta}_{k,\delta}\right)\right)$
$\leq$ $\exp\left(\dfrac{n}{2k}\ln\left(2((p_{k}^{\max}-\bar{\eta}_{k,\delta})(1-p_{k}^{\max}+\bar{\eta}_{k,\delta}))^{1/5}\right)\right)$
$\leq$ $\left(2(p_{k}^{\max}-\bar{\eta}_{k,\delta})^{1/5}\right)^{n/(2k)}$
$=$ $\left(2^{5}(p_{k}^{\max}-\bar{\eta}_{k,\delta})\right)^{n/(10k)},$
and similarly
$\exp\left(-\dfrac{n}{2k}\cdot D_{\operatorname{KL}}\left(p_{k}^{\max}-\dfrac{\bar{\eta}_{k,\delta}}{4}\Big{\|}p_{k}^{\max}\right)\right)$
$\leq$ $\exp\left(\dfrac{n}{2k}\ln\left(2(p_{k}^{\max}(1-p_{k}^{\max}))^{1/5}\right)\right)$
$=$ $\left(2(p_{k}^{\max}(1-p_{k}^{\max}))^{1/5}\right)^{n/(2k)}$
$\leq$ $\left(2(1-p_{k}^{\max})^{1/5}\right)^{n/(2k)}$
$=$ $\left(2^{5}(1-p_{k}^{\max})\right)^{n/(10k)}.$
On the other hand, by Lemma 3 both $D_{\operatorname{KL}}\left(p_{k}^{\max}-3\bar{\eta}_{k,\delta}/4\|p_{k}^{\max}-\bar{\eta}_{k,\delta}\right)$ and $D_{\operatorname{KL}}\left(p_{k}^{\max}-\bar{\eta}_{k,\delta}/4\|p_{k}^{\max}\right)$ are bounded below by $\bar{\eta}_{k,\delta}^{2}/8$, therefore
$\dfrac{\bar{\eta}_{k,\delta}^{2}}{\min(p_{k}^{\max},1-p_{k}^{\max})+3\bar{\eta}_{k,\delta}/4}\geq\dfrac{(4/5)^{2}}{\min(1,1/5)+3/4}\geq\dfrac{16}{25},$
and hence
$\exp\left(-\dfrac{B}{24}\cdot\dfrac{\bar{\eta}_{k,\delta}^{2}}{\min\left(p_{k}^{\max},1-p_{k}^{\max}\right)+3\bar{\eta}_{k,\delta}/4}\right)\leq\exp\left(-\dfrac{B}{75/2}\right).$
The first desired bound then follows by setting $C_{1},C_{2},C_{3}$ to be the appropriate constants. This completes the proof of Theorem 3. $\Box$
Proof of Theorem 1. Algorithm 1 is a special case of Algorithm 3 with the learning algorithm (23) that outputs a singleton, therefore the results of Theorem 3 automatically apply. Since $\mathcal{E}_{k,\delta}=\sum_{\theta\in\Theta\backslash\Theta^{\delta}}p_{k}(\theta)\geq\max_{\theta\in\Theta\backslash\Theta^{\delta}}p_{k}(\theta)$, it holds that $\eta_{k,\delta}\leq\bar{\eta}_{k,\delta}$. When $\eta_{k,\delta}>0$ we also have $\bar{\eta}_{k,\delta}>0$, therefore (8) follows from setting $\eta$ to be $\eta_{k,\delta}$ in (30), and (9) follows from upper bounding $\max_{\theta\in\Theta\backslash\Theta^{\delta}}p_{k}(\theta)$ with $\mathcal{E}_{k,\delta}$ in (31). $\Box$
C.3 Proof for Theorem 2
We first present two lemmas to be used in the main proof. The first lemma characterizes the exponentially improving quality of the solution set retrieved in Phase I:
Lemma 6 (Quality of retrieved solutions in Algorithm 2)
$\mathbb{P}\left(\mathcal{S}\cap\Theta^{\delta}=\emptyset\right)\leq\min\left(e^{-(1-\mathcal{E}_{k,\delta})/C_{4}},C_{5}\mathcal{E}_{k,\delta}\right)^{\frac{n}{C_{6}k}}+\exp\left(-\dfrac{B_{1}}{C_{7}}(1-\mathcal{E}_{k,\delta})\right),$ (37)
@@ C.3 Proof for Theorem 2
Proof of Lemma 6. Let $(z_{1}^{*},\ldots,z_{k}^{*})$ be a random subsample and $\mathbb{P}_{*}$ be the probability with respect to the subsampling randomness conditioned on the data and the algorithmic randomness. Consider the two probabilities
@@
$\leq$ $\mathbb{P}\left(\mathbb{P}_{*}\left(\mathcal{A}(z_{1}^{*},\ldots,z_{k}^{*})\in\Theta^{\delta}\right)<\frac{1-\mathcal{E}_{k,\delta}}{e}\right)+\left(1-\frac{1-\mathcal{E}_{k,\delta}}{e}\right)^{B_{1}}$
where $e$ is the base of the natural logarithm. Applying Lemma 2 with $\kappa(z_{1},\ldots,z_{k};\omega):=\mathbbm{1}\left(\mathcal{A}(z_{1},\ldots,z_{k};\omega)\in\Theta^{\delta}\right)$ gives
@@
Further applying the bound (17) from Lemma 3 to the KL divergence on the right-hand side leads to
$D_{\operatorname{KL}}\left(\frac{1-\mathcal{E}_{k,\delta}}{e}\Big{\|}1-\mathcal{E}_{k,\delta}\right)\geq\frac{1-\mathcal{E}_{k,\delta}}{e}\ln\frac{1}{e}+1-\mathcal{E}_{k,\delta}-\frac{1-\mathcal{E}_{k,\delta}}{e}=\left(1-\frac{2}{e}\right)(1-\mathcal{E}_{k,\delta})$
@@
$D_{\operatorname{KL}}\left(\frac{1-\mathcal{E}_{k,\delta}}{e}\Big{\|}1-\mathcal{E}_{k,\delta}\right)$
$=$ $D_{\operatorname{KL}}\left(1-\frac{1-\mathcal{E}_{k,\delta}}{e}\Big{\|}\mathcal{E}_{k,\delta}\right)$

### v4

@@ Appendix D Proofs for Technical Lemmas
This completes the proof. ∎
D.2 Proof of Lemma C.3
We first consider the direction $U-\mathbb{E}\left[\kappa\right]\geq\epsilon$. Let
@@
where we apply the Markov inequality in $(i)$, step $(ii)$ is due to Lemma C.2, and step $(iii)$ uses Jensen’s inequality and the convexity of the exponential function. Due to independence, $\tilde{\kappa}$ can be viewed as the sample average of $\hat{n}$ i.i.d. Bernoulli random variables, i.e., $\tilde{\kappa}\sim\frac{1}{\hat{n}}\sum_{i=1}^{\hat{n}}\operatorname{Bernoulli}\left(\mathbb{E}\left[\kappa\right]\right)$. Hence, we have that
@@
Since $D_{\operatorname{KL}}(p\|q)=D_{\operatorname{KL}}(1-p\|1-q)$, the bound for the reverse side $U-\mathbb{E}\left[\kappa\right]\leq-\epsilon$ then follows by applying the first bound to the flipped binary kernel $1-\kappa$ and $1-U$. This completes the proof of Lemma C.3. ∎
D.3 Proof of Lemma C.4
To show (13), some basic calculus shows that for any fixed $q$, the function $g(p):=(1-p)\ln\frac{1-p}{1-q}$ is convex in $p$, and we have that
@@
This completes the proof of Lemma C.4. ∎
D.4 Proof of Lemma C.9
By Definition C.5, we observe the following equivalence
@@
$\mathbb{P}\left({\hat{p}}_{k}(\theta)\geq{\hat{p}}_{k}\left(\theta_{k}^{\max}\right)-\epsilon\right)$$\leq$ ($\mathbb{P}\left(\left\{{\hat{p}}_{k}(\theta)\geq p_{k}(\theta_{k}^{\max})-\dfrac{\delta+\epsilon}{2}\right\}\cap\left\{{\hat{p}}_{k}\left(\theta_{k}^{\max}\right)\leq p_{k}(\theta_{k}^{\max})-\dfrac{\delta-\epsilon}{2}\right\}\right)$ (50)
@@
Note that, with $\kappa(z_{1},\ldots,z_{k};\omega):=\mathbf{1}\left(\theta\in\mathbb{A}(z_{1},\ldots,z_{k};\omega)\right)$, the probability ${\hat{p}}_{k}(\theta)$ can be viewed as a U-statistic with the kernel $\kappa^{*}(z_{1},\ldots,z_{k}):=\mathbb{E}\left[\kappa(z_{1},\ldots,z_{k};\omega)|z_{1},\ldots,z_{k}\right]$. A similar representation holds for ${\hat{p}}_{k}\left(\theta_{k}^{\max}\right)$ as well. Therefore, we can apply Lemma C.3 to conclude that
@@
which completes the proof of Lemma C.9. ∎
D.5 Proof of Lemma C.10
We observe that $\bar{p}_{k}(\theta)$ is an conditionally unbiased estimator for $\hat{p}_{k}(\theta)$, i.e., $\mathbb{E}_{*}\left[\bar{p}_{k}(\theta)\right]={\hat{p}}_{k}(\theta)$. We can express the difference between $\bar{p}_{k}(\theta)$ and $\bar{p}_{k}(\theta_{k}^{\max})$ as the sample average
@@
$\mathbb{P}_{*}\left(\bar{p}_{k}(\theta)-\bar{p}_{k}(\hat{\theta}_{k}^{\max})-(\hat{p}_{k}(\theta)-\hat{p}_{k}(\theta_{k}^{\max}))\geq t\right)\leq$ $\exp\left(-B\cdot\dfrac{t^{2}}{2\mathrm{Var}_{*}(\mathbbm{1}_{\theta}^{*}-\mathbbm{1}_{\theta_{k}^{\max}}^{*})+4/3\cdot t}\right).$ (51)
@@
$\mathrm{Var}_{*}(\mathbbm{1}_{\theta}^{*}-\mathbbm{1}_{\theta_{k}^{\max}}^{*})$$\leq$ $\mathbb{E}_{*}\left[(\mathbbm{1}_{\theta}^{*}-\mathbbm{1}_{\theta_{k}^{\max}}^{*})^{2}\right]\leq\hat{p}_{k}(\theta)+\hat{p}_{k}(\theta_{k}^{\max})\leq 2\hat{p}_{k}(\theta_{k}^{\max}),$
@@
we have $\mathrm{Var}_{*}(\mathbbm{1}_{\theta}^{*}-\mathbbm{1}_{\theta_{k}^{\max}}^{*})\leq 2\min\left\{\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta)\right\}$. Substituting this bound to (51) and taking $t=\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta)$ lead to
$\mathbb{P}_{*}\left(\bar{p}_{k}(\theta)-\bar{p}_{k}(\hat{\theta}_{k}^{\max})\geq 0\right)$$\leq$ $\exp\left(-B\cdot\dfrac{(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))^{2}}{4\min\left\{\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta)\right\}+4/3\cdot(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))}\right)$
$\leq$ $\exp\left(-B\cdot\dfrac{(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))^{2}}{4\min\left\{\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta_{k}^{\max})\right\}+16/3\cdot(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))}\right)$
$\leq$ $\exp\left(-\dfrac{B}{6}\cdot\dfrac{(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))^{2}}{\min\left\{\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta_{k}^{\max})\right\}+\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta)}\right).$
@@
$\leq\sum_{\theta\in\Theta\backslash\widehat{\mathcal{P}}^{\epsilon}_{k}}\exp\left(-\dfrac{B}{6}\cdot\dfrac{(\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta))^{2}}{\min\left\{\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta_{k}^{\max})\right\}+\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta)}\right).$
@@
Note that the function $x^{2}/(\min\left\{\hat{p}_{k}(\theta_{k}^{\max}),1-\hat{p}_{k}(\theta_{k}^{\max})\right\}+x)$ in $x\in[0,1]$ is monotonically increasing and that $\hat{p}_{k}(\theta_{k}^{\max})-\hat{p}_{k}(\theta)>\epsilon$ for all $\theta\in\Theta\backslash\widehat{\mathcal{P}}^{\epsilon}_{k}$. Therefore, we can further bound the probability as
@@
$\mathbb{P}_{*}\left(\hat{\theta}_{n}\notin\widehat{\mathcal{P}}^{\epsilon}_{k}\right)$$\leq$ $\left|\Theta\backslash\widehat{\mathcal{P}}^{\epsilon}_{k}\right|\cdot\exp\left(-\dfrac{B}{6}\cdot\dfrac{\epsilon^{2}}{\min\left\{\hat{p}_{k}^{\max},1-\hat{p}_{k}^{\max}\right\}+\epsilon}\right).$
@@
Noting that $\left|\Theta\backslash\widehat{\mathcal{P}}^{\epsilon}_{k}\right|\leq\left|\Theta\right|$ completes the proof of Lemma C.10. ∎
D.6 Proof of Lemma C.12
@@
Let $(z_{1}^{*},\ldots,z_{k}^{*})$ be a random subsample and $\mathbb{P}_{*}$ be the probability with respect to the subsampling randomness conditioned on the data and the algorithmic randomness. Consider the two probabilities
@@
$\leq$ $\mathbb{P}\left(\mathbb{P}_{*}\left(\mathcal{A}(z_{1}^{*},\ldots,z_{k}^{*})\in\Theta^{\delta}\right)<\frac{1-\mathcal{E}_{k,\delta}}{e}\right)+\left(1-\frac{1-\mathcal{E}_{k,\delta}}{e}\right)^{B_{1}},$
where $e$ is the base of the natural logarithm. Applying Lemma C.3 with $\kappa(z_{1},\ldots,z_{k};\omega):=\mathbbm{1}\left(\mathcal{A}(z_{1},\ldots,z_{k};\omega)\in\Theta^{\delta}\right)$ gives
@@
Further applying the bound (13) from Lemma C.4 to the KL divergence on the right-hand side leads to
$D_{\operatorname{KL}}\left(\frac{1-\mathcal{E}_{k,\delta}}{e}\Big{\|}1-\mathcal{E}_{k,\delta}\right)\geq\frac{1-\mathcal{E}_{k,\delta}}{e}\ln\frac{1}{e}+1-\mathcal{E}_{k,\delta}-\frac{1-\mathcal{E}_{k,\delta}}{e}=\left(1-\frac{2}{e}\right)(1-\mathcal{E}_{k,\delta}),$
@@
$D_{\operatorname{KL}}\left(\frac{1-\mathcal{E}_{k,\delta}}{e}\Big{\|}1-\mathcal{E}_{k,\delta}\right)$ $=$ $D_{\operatorname{KL}}\left(1-\frac{1-\mathcal{E}_{k,\delta}}{e}\Big{\|}\mathcal{E}_{k,\delta}\right)$